find the number of molecules of oxygen required to convert 6.60 dm^3 of SO2 measured at s.t.p to SO3


1
Expert's answer
2021-02-25T03:44:34-0500

2SO2 + O2 "\to" 2SO3

6.60dm3=6.6L

n(SO2)=6.6L/22.4L=0.2946mol

n(O2)=0.2946mol/2=0.1473mol

N(O2)=0.1473×6.02×1023 = 0.887×1023

= 8.87 × 1022